# SpeakerLM: End-to-End Versatile Speaker Diarization and Recognition with Multimodal Large Language Models

### v1

@@ Introduction
Speaker Diarization (SD) partitions an audio into homogeneous segments according to speaker identities, effectively answering “who spoke when” (Tranter and Reynolds 2006; Anguera et al. 2012; Park et al. 2022). In real-world scenarios such as multi-speaker meetings, knowing “who spoke when” is insufficient. Associating speaker labels with speech transcripts is more meaningful, as it addresses the more comprehensive question of “who spoke when and what” (Gao et al. 2025). We refer to this task as Speaker Diarization and Recognition (SDR), which jointly performs SD and Automatic Speech Transcription (ASR) for enriched conversational understanding.
@@
Conventional SDR systems typically pipeline an SD module and an ASR module (Raj et al. 2021; Yu et al. 2022; Cornell et al. 2024; Boeddeker et al. 2024). The SD module segments the audio stream into speaker-homogeneous regions, which are then aligned with the ASR outputs, and the speech content for each identified speaker is transcribed. However, such cascaded systems suffer from several limitations. First, error propagation is a major concern. Inaccuracies in SD, such as incorrect speaker boundaries or label assignments, are passed on to the ASR module, leading to degraded transcription quality and incorrect speaker-attributed text (Cornell et al. 2023). In addition, conventional SD systems struggle to handle overlapping speech, which is common in real-world conversations, resulting in limited performance (O’Shaughnessy 2025). Furthermore, the cascaded system lacks joint optimization as ASR and SD modules are usually trained independently with different datasets and frameworks. The disjoint training paradigm prevents the SDR system from fully leveraging the synergy between SD and ASR tasks, limiting the overall SDR performance.
@@
Two main paradigms have been attempted to address these limitations of cascaded SDR systems. One paradigm, namely Speaker-Attributed ASR (SA-ASR), jointly trains SD and ASR. As a representative end-to-end SA-ASR system, SA-Transformer (Liang et al. 2023) incorporates an ASR encoder and a speaker encoder, and achieves competitive SDR results. However, SA-ASR systems can only operate under the condition that pre-extracted speaker embeddings are available. Another paradigm uses Large Language Models (LLMs) (Wei et al. 2022; Chang et al. 2024) for post-processing. Due to the disjoint training issue in cascaded SDR systems, temporal mismatches between the two modules may lead to word-level diarization errors (Wang et al. 2024). Prior works (Park et al. 2024; Efstathiadis, Yadav, and Abbas 2025) explore the strong generalization and reasoning abilities of LLMs to post-process the outputs of a cascaded SDR system, refining speaker attributions and correcting transcription inconsistencies. While such post-hoc adjustments may improve the overall SDR performance, they are inherently limited by the quality of the initial SD and ASR outputs, and cannot fully resolve issues arising from upstream error propagation or misalignment between modalities. Therefore, how to leverage LLMs not just for post-processing but as core components in end-to-end SDR systems remains an open challenge.
@@
Related Work
Speaker Diarization
Traditional speaker diarization frameworks typically consist of multiple components, including Voice Activity Detection (VAD) to identify speech segments, speaker embedding extraction to represent speaker characteristics, and clustering to group speaker embeddings for identifying speakers. 3D-Speaker (Chen et al. 2025b) is a representative traditional SD system following this pipeline, and achieves competitive performance across multiple public benchmarks.
Conventional SD systems struggle with overlapping speech, because VAD is designed under the assumption of single-speaker segments. One promising solution to address this issue is to replace the VAD module with a neural speaker activity detection module, which predicts active speakers in each time frame (Fujita et al. 2019; Kinoshita, Delcroix, and Tawara 2021b, a). Along this direction, Pyannote (Bredin et al. 2020; Bredin 2023) achieves strong performance on different datasets. Based on Pyannote, Diarizen (Han et al. 2025a, b) incorporates representations from the pretrained WavLM model (Chen et al. 2022) into the neural speaker activity detection module, outperforming Pyannote on various datasets by providing richer acoustic information.
@@
Speaker Diarization and Recognition
Figure 1 compares three system architectures for SDR. Figure 1 (a) depicts a conventional SDR system that pipelines SD and ASR modules (Cornell et al. 2024), denoted as SD+ASR. This pipeline has several limitations, such as error propagation and lack of joint optimization. To address these issues, one direction uses LLMs for post-processing (Efstathiadis, Yadav, and Abbas 2025), and another direction jointly models SD and ASR tasks within a single unified model in an end-to-end manner. We denote these two methods by SD+ASR+LLM and E2E-SDR, respectively.
For E2E-SDR, existing studies typically incorporate pre-extracted speaker embeddings for all speakers. This framework, also referred to as SA-ASR (Liang et al. 2023), assumes that the provided speaker embeddings exactly match the speakers in the ground truth. Consequently, it struggles to handle mismatched scenarios such as unregistered speakers or over-registered cases. Our work follows the E2E-SDR paradigm, but is different from the prior SA-ASR systems. Specifically, we introduce a more flexible speaker registration mechanism during training, enabling the model to handle speaker registration inconsistencies effectively. Compared with the cascaded frameworks, our unified modeling approach allows for exploring the synergy between SD and ASR tasks and joint optimization, leading to better alignment between speaker and textual information.
Multimodal Large Language Model
Recently, integrating audio and text with MLLMs has led to notable advances in auditory comprehension and reasoning. Representative models, such as MinMo (Chen et al. 2025a), Qwen2-Audio (Chu et al. 2024) and Kimi-Audio (Ding et al. 2025), have demonstrated strong capabilities in audio-text understanding and generation tasks. However, most existing MLLMs primarily focus on single-speaker scenarios, where they exhibit strong performance in ASR tasks but leave SD tasks unexplored. More recently, the Multi-Talker LLM (MT-LLM) has been proposed to handle multi-speaker ASR (Meng et al. 2025). However, MT-LLM is limited to detecting speaker change points and do not explicitly tackle speaker attribution within multi-speaker transcripts.
@@
SpeakerLM
Model Architecture
@@
Audio Encoder and Projector. The audio encoder is initialized with the pre-trained SenseVoice-large encoder (An et al. 2024), which offers robust audio representation capabilities and provides strong performance across various audio understanding tasks such as multilingual speech recognition and audio event detection. For the audio projector, we use a randomly initialized two-layer Transformer (Vaswani et al. 2017), followed by a convolutional neural network layer for dimensional alignment.
Embedding Extractor and Projector. A pre-trained speaker embedding model is utilized to extract speaker embeddings, delivering robust and discriminative representations that are essential for precise speaker identification and attribution. Specifically, we use an open-source speaker embedding model, i.e., ERes2NetV2 (Chen et al. 2024) for embedding extraction, which has achieved SOTA performance on multiple speaker verification benchmarks. A single linear layer projector is applied to the extracted embedding for dimensional alignment.
Large Language Model. We use the pre-trained Qwen2.5-7B-Instruct model (Team 2024) as the backbone text LLM to benefit from its strong instruction-following and general language understanding capabilities, which allows SpeakerLM to effectively handle complex multi-speaker SDR tasks with different levels of information.
Flexible Speaker Registration Mechanism
@@
Multi-Stage Training Strategy
@@
Stage 1. To enhance the ASR performance, we first train SpeakerLM using only ASR data, leading to an ASR model, denoted by SpeakerLM-ASR. To efficiently adapt the LLM while preserving its pre-trained language capabilities, we apply a Low-Rank Adaptation (LoRA) (Hu et al. 2022) strategy. Following the ASR training setup of MinMo (Chen et al. 2025a), we sample 600,000 hours of audio from publicly available ASR datasets for training. In this stage, the speaker embedding extractor and projector are not included in the model. In the subsequent three stages, these modules are incorporated into the whole architecture.
@@
Data Composition
Real Data Collection
@@
Specifically, the public Alimeeting (Liang et al. 2023) and AISHELL4 (Fu et al. 2021) datasets are collections of multi-speaker conversation recordings in meeting rooms, including both far-field overlapping audio clips and near-field recordings for each speaker. Following previous works (Han et al. 2025a; Chen et al. 2025b), we use the first channel of the far-field signals for training and evaluation. The in-house data contains multi-speaker audio samples recorded in professional studios. The AISHELL5 dataset (Dai et al. 2025) was recorded using far-field microphones in a hybrid electric car. The dataset includes diverse external noise sources such as wind and tire, along with internal noises from music and air conditioning, making it a more challenging setting than standard meeting room or studio environments. The dataset has two official test set splits for ASR and SDR tasks respectively. We use the first channel of the far-field audio from the SDR test split as the test set in this work, denoted by AISHELL5-Eval. Notably, AISHELL5-Eval presents a challenging acoustic condition that is significantly different from the domains of the training data, hence it is primarily used to evaluate the out-of-domain SDR performance (generalizability) of the proposed SpeakerLM.
@@
Simulated Data Creation
As described in Section Multi-Stage Training Strategy, we apply a four-stage training strategy to train SpeakerLM. In the second stage, we use simulated data for quickly training the projectors and audio-text alignment. Specifically, we use the near-field speech samples from AliMeeting-Train, AISHELL2 (Du et al. 2018), Librispeech (Panayotov et al. 2015) and In-House-Train for SDR data simulation. The near-field speech samples are randomly mixed to simulate multi-speaker conversation segments, with each segment lasting 50 seconds and containing 2 to 4 speakers. Following previous work (Landini et al. 2022), we randomly add real-world noises and reverberations to the clean speech samples, with signal-to-noise ratios (SNRs) uniformly sampled from [10dB, 20dB]. In total, the simulated data comprise 5,000 hours audio samples for training and 5.6 hours for evaluation; the test split is denoted by Simulation-Test.
Experiments
Implementation Details
@@
We use the AdamW (Zhou et al. 2024) optimizer for optimization. The learning rate is initially set to 1e-5, linearly increased to 5e-5 during the warm-up phase, and then gradually decayed following a cosine schedule. We use a dynamic batching strategy, adjusting the number of samples per batch based on their length, with a maximum token limit of 6K to balance efficiency and memory usage. Models are trained on 4 NVIDIA A800 GPUs for 1M steps at each stage, with a validation step of 10K. The training pseudocode and the specific prompts used for LLMs are presented in Appendix.
@@
Evaluation Setups
Evaluation Metrics. We evaluate the SDR performance on the commonly used publicly available benchmarks, namely, the in-domain AliMeeting-Eval and AISHELL4-Eval, and the out-of-domain AISHELL5-Eval. Following previous works (Liang et al. 2023; Wang et al. 2024; Dai et al. 2025), we use Character Error Rate (CER), concatenated minimum permutation CER (cpCER), $\Delta\textrm{cp}$, speaker-attributed CER (saCER), and $\Delta\textrm{sa}$ as evaluation metrics.
@@
For the SD+ASR paradigm, we use Paraformer-large (Gao et al. 2022) as the ASR module, which was pre-trained on 60K hours of audio data and achieves SOTA performance on various Mandarin ASR benchmarks. For the SD module, we apply four SOTA open-source SD toolkits, including 3D-Speaker (Chen et al. 2025b), Pyannote 3.1 (Bredin 2023), Diarizen-base (Han et al. 2025a), and Diarizen-large (Han et al. 2025b). Among the four SD models, 3D-Speaker and Diarizen-large have achieved SOTA performance on different benchmarks, while Pyannote 3.1 and Diarizen-base demonstrate competitive performance. By combining the four SD systems with Paraformer-large, we obtain four SD+ASR baseline systems, namely, 3D-Speaker+Para, Pyannote+Para, Diarizen-base+Para, and Diarizen-large+Para.
SD+ASR+LLM refers to a pipeline where a text LLM is used to correct the outputs from the SD+ASR system. For SD+ASR+LLM baselines, prior works using a text LLM for SD+ASR correction are mainly based on English datasets (Wang et al. 2024; Park et al. 2024; Efstathiadis, Yadav, and Abbas 2025). Following the paradigm, we adapt Qwen2.5-7B-Instruct (Team 2024) and ChatGPT4.5 (Singh 2025) for correction. We first evaluate their zero-shot performance (z.s.) using prompts following prior work (Efstathiadis, Yadav, and Abbas 2025), then fine-tune Qwen2.5-7B-Instruct (f.t.) on our training datasets for SD+ASR correction.
Performance without Speaker Registration
@@
For SD+ASR+LLM baselines pipelined with five models, we incorporate a pre-trained text LLM to correct the mis-assigned speaker labels in the results from the best SD+ASR system (i.e., Diarizen-large+Para). Results show that both Qwen2.5-7B-Instruct and ChatGPT-4.5 perform poorly in zero-shot SDR correction, even degrading the overall performance. This finding is consistent with observations in prior works (Park et al. 2024; Efstathiadis, Yadav, and Abbas 2025). This is mainly due to hallucination of LLMs: despite that we explicitly instruct the LLM to only modify speaker labels, the LLM often alters the content of the speakers’ utterances, leading to a significant increase in CER. Finetuning Qwen2.5-7B-Instruct substantially mitigates the hallucination problem, leading to slight SDR performance improvement over Diarizen-large+Para.
Different from the cascaded architectures of SD+ASR and SD+ASR+LLM, the proposed SpeakerLM is a unified E2E-SDR model that jointly performs speaker diarization and recognition. To explore how much SpeakerLM benefits from the ASR training stage, we first report the CER of SpeakerLM-ASR, which is trained on the ASR data mentioned in Section Multi-Stage Training Strategy. Compared with Paraformer-large, SpeakerLM-ASR is trained on a larger ASR dataset and thus achieves lower CER across all test sets, which naturally contributes to a lower cpCER for SpeakerLM. However, it should be noted that $\Delta\textrm{cp}$ is not affected by the model’s ASR performance, which serves as a reliable metric for evaluating the speaker attribution performance.
@@
Impact of Multi-Stage Training. Figure 4 illustrates the performance of SpeakerLM across the four different training stages (Section Multi-Stage Training Strategy). On both Alimeeting-Eval and Simulation-Test, the performance of SpeakerLM improves progressively, demonstrating the effectiveness of the proposed multi-stage training strategy. However, on AISHELL4-Eval and AISHELL5-Eval, CERs at Stage 2 are higher than those at Stage 1. This is because Stage 2 relies on simulated data, while the simulation process does not incorporate any audio data from these two datasets, leading to a domain mismatch. With four training stages, SpeakerLM demonstrates strong generalizability on the out-of-domain test set, i.e, AISHELL5-Eval. This suggests that the subsequent training stages (Stage 3 and Stage 4), which involve fine-tuning on more realistic and diverse meeting-style data, are crucial for mitigating domain mismatch and enhancing the model’s robustness across different evaluation scenarios. In addition, the ASR performance on simulated data is consistently better than that on real data, which is attributed to the simplicity of the simulated data. Nevertheless, we believe the simulation process remains meaningful, as it closely resembles office or workplace scenarios where a voice assistant is used to record meetings.
@@
Performance with Speaker Registration

### v2

@@ 1 Introduction
Speaker Diarization (SD) partitions an audio into homogeneous segments according to speaker identities, effectively answering “who spoke when” (sd1_2006; sd2_2012; sd3_2022). In real-world scenarios such as multi-speaker meetings, knowing “who spoke when” is insufficient. Associating speaker labels with speech transcripts is more meaningful, as it addresses the more comprehensive question of “who spoke when and what” (misp2025). We refer to this task as Speaker Diarization and Recognition (SDR), which jointly performs SD and Automatic Speech Transcription (ASR) for enriched conversational understanding.
@@
Conventional SDR systems typically pipeline an SD module and an ASR module (sd+asr2_2021; sd_asr3_2022; sd+asr1_2024; sd_asr4_2024). The SD module segments the audio stream into speaker-homogeneous regions, which are then aligned with the ASR outputs, and the speech content for each identified speaker is transcribed. However, such cascaded systems suffer from several limitations. First, error propagation is a major concern. Inaccuracies in SD, such as incorrect speaker boundaries or label assignments, are passed on to the ASR module, leading to degraded transcription quality and incorrect speaker-attributed text (sd_asr5_2023). In addition, conventional SD systems struggle to handle overlapping speech, which is common in real-world conversations, resulting in limited performance (sd4_2025). Furthermore, the cascaded system lacks joint optimization as ASR and SD modules are usually trained independently with different datasets and frameworks. The disjoint training paradigm prevents the SDR system from fully leveraging the synergy between SD and ASR tasks, limiting the overall SDR performance.
@@
Two main paradigms have been attempted to address these limitations of cascaded SDR systems. One paradigm, namely Speaker-Attributed ASR (SA-ASR), jointly trains SD and ASR. As a representative end-to-end SA-ASR system, SA-Transformer (alimeeting_2023) incorporates an ASR encoder and a speaker encoder, and achieves competitive SDR results. However, SA-ASR systems can only operate under the condition that pre-extracted speaker embeddings are available. Another paradigm uses Large Language Models (LLMs) (llm2_2022; llm1_2024) for post-processing. Due to the disjoint training issue in cascaded SDR systems, temporal mismatches between the two modules may lead to word-level diarization errors (diarizationlm_2024). Prior works (sd_llm_3_2024; sd_llm_2_2025) explore the strong generalization and reasoning abilities of LLMs to post-process the outputs of a cascaded SDR system, refining speaker attributions and correcting transcription inconsistencies. While such post-hoc adjustments may improve the overall SDR performance, they are inherently limited by the quality of the initial SD and ASR outputs, and cannot fully resolve issues arising from upstream error propagation or misalignment between modalities. Therefore, how to leverage LLMs not just for post-processing but as core components in end-to-end SDR systems remains an open challenge.
@@
2 Related Work
2.1 Speaker Diarization
Traditional speaker diarization frameworks typically consist of multiple components, including Voice Activity Detection (VAD) to identify speech segments, speaker embedding extraction to represent speaker characteristics, and clustering to group speaker embeddings for identifying speakers. 3D-Speaker (3d_speaker_2025) is a representative traditional SD system following this pipeline, and achieves competitive performance across multiple public benchmarks.
Conventional SD systems struggle with overlapping speech, because VAD is designed under the assumption of single-speaker segments. One promising solution to address this issue is to replace the VAD module with a neural speaker activity detection module, which predicts active speakers in each time frame (e2e_sd_2019; e2e_vc_sd_1_2021; e2e_vc_sd_2_2021). Along this direction, Pyannote (pyannote_1_2020; pyannote_2_2023) achieves strong performance on different datasets. Based on Pyannote, Diarizen (diarizen_1_2025; diarizen_2_2025) incorporates representations from the pretrained WavLM model (wavlm_2022) into the neural speaker activity detection module, outperforming Pyannote on various datasets by providing richer acoustic information.
@@ 2.1 Speaker Diarization
2.2 Speaker Diarization and Recognition
Figure 1 compares three system architectures for SDR. Figure 1 (a) depicts a conventional SDR system that pipelines SD and ASR modules (sd+asr1_2024), denoted as SD+ASR. This pipeline has several limitations, such as error propagation and lack of joint optimization. To address these issues, one direction uses LLMs for post-processing (sd_llm_2_2025), and another direction jointly models SD and ASR tasks within a single unified model in an end-to-end manner. We denote these two methods by SD+ASR+LLM and E2E-SDR, respectively.
For E2E-SDR, existing studies typically incorporate pre-extracted speaker embeddings for all speakers. This framework, also referred to as SA-ASR (alimeeting_2023), assumes that the provided speaker embeddings exactly match the speakers in the ground truth. Consequently, it struggles to handle mismatched scenarios such as unregistered speakers or over-registered cases. Our work follows the E2E-SDR paradigm, but is different from the prior SA-ASR systems. Specifically, we introduce a more flexible speaker registration mechanism during training, enabling the model to handle speaker registration inconsistencies effectively. Compared with the cascaded frameworks, our unified modeling approach allows for exploring the synergy between SD and ASR tasks and joint optimization, leading to better alignment between speaker and textual information.
2.3 Multimodal Large Language Model
Recently, integrating audio and text with MLLMs has led to notable advances in auditory comprehension and reasoning. Representative models, such as MinMo (minmo_2025), Qwen2-Audio (qwen2-audio_2024) and Kimi-Audio (kimi-audio_2025), have demonstrated strong capabilities in audio-text understanding and generation tasks. However, most existing MLLMs primarily focus on single-speaker scenarios, where they exhibit strong performance in ASR tasks but leave SD tasks unexplored. More recently, the Multi-Talker LLM (MT-LLM) has been proposed to handle multi-speaker ASR (mt_llm_2025). However, MT-LLM is limited to detecting speaker change points and do not explicitly tackle speaker attribution within multi-speaker transcripts.
@@ 2.3 Multimodal Large Language Model
3 SpeakerLM
3.1 Model Architecture
@@
Audio Encoder and Projector. The audio encoder is initialized with the pre-trained SenseVoice-large encoder (sensevoice_2024), which offers robust audio representation capabilities and provides strong performance across various audio understanding tasks such as multilingual speech recognition and audio event detection. For the audio projector, we use a randomly initialized two-layer Transformer (transformer_2017), followed by a convolutional neural network layer for dimensional alignment.
Embedding Extractor and Projector. A pre-trained speaker embedding model is utilized to extract speaker embeddings, delivering robust and discriminative representations that are essential for precise speaker identification and attribution. Specifically, we use an open-source speaker embedding model ERes2NetV2 (eres2netv2_2024) for embedding extraction, which has achieved SOTA performance on multiple speaker verification benchmarks. A single linear layer projector is applied to the extracted embedding for dimensional alignment.
Large Language Model. We use the pre-trained Qwen2.5-7B-Instruct model (qwen2.5_2024) as the backbone text LLM to benefit from its strong instruction-following and general language understanding capabilities, which allows SpeakerLM to effectively handle complex multi-speaker SDR tasks with different levels of information.
3.2 Flexible Speaker Registration Mechanism
@@
3.3 Multi-Stage Training Strategy
@@
Stage 1. To enhance the ASR performance, we first train SpeakerLM using only ASR data, leading to an ASR model, denoted by SpeakerLM-ASR. To efficiently adapt the LLM while preserving its pre-trained language capabilities, we apply a Low-Rank Adaptation (LoRA) (lora_2022) strategy. Following the ASR training setup of MinMo (minmo_2025), we sample 600,000 hours of audio from publicly available ASR datasets for training. In this stage, the speaker embedding extractor and projector are not included in the model. In the subsequent three stages, these modules are incorporated into the whole architecture.
@@
4 Data Composition
4.1 Real Data Collection
@@
Specifically, the public Alimeeting (alimeeting_2023) and AISHELL4 (aishell4_2021) datasets are collections of multi-speaker conversation recordings in meeting rooms, including both far-field overlapping audio clips and near-field recordings for each speaker. Following previous works (diarizen_1_2025; 3d_speaker_2025), we use the first channel of the far-field signals for training and evaluation. The in-house data contains multi-speaker audio samples recorded in professional studios. The AISHELL5 dataset (aishell5_2025) was recorded using far-field microphones in a hybrid electric car. The dataset includes diverse external noise sources such as wind and tire, along with internal noises from music and air conditioning, making it a more challenging setting than standard meeting room or studio environments. The dataset has two official test set splits for ASR and SDR tasks respectively. We use the first channel of the far-field audio from the SDR test split as the test set in this work, denoted by AISHELL5-Eval. Notably, AISHELL5-Eval presents a challenging acoustic condition that is significantly different from the domains of the training data, hence it is primarily used to evaluate the out-of-domain SDR performance (generalizability) of the proposed SpeakerLM.
@@
4.2 Simulated Data Creation
As described in Section 3.3, we apply a four-stage training strategy to train SpeakerLM. In the second stage, we use simulated data for quickly training the projectors and audio-text alignment. Specifically, we use the near-field speech samples from AliMeeting-Train, AISHELL2 (aishell2_2018), Librispeech (librispeech_2015) and In-House-Train for SDR data simulation. The near-field speech samples are randomly mixed to simulate multi-speaker conversation segments, with each segment lasting 50 seconds and containing 2 to 4 speakers. Following previous work (simulated_2022), we randomly add real-world noises and reverberations to the clean speech samples, with signal-to-noise ratios (SNRs) uniformly sampled from [10dB, 20dB]. In total, the simulated data comprise 5,000 hours audio samples for training and 5.6 hours for evaluation; the test split is denoted by Simulation-Test.
5 Experiments
5.1 Implementation Details
@@
We use the AdamW (adamw_2024) optimizer for optimization. The learning rate is initially set to 1e-5, linearly increased to 5e-5 during the warm-up phase, and then gradually decayed following a cosine schedule. We use a dynamic batching strategy, adjusting the number of samples per batch based on their length, with a maximum token limit of 6K to balance efficiency and memory usage. Models are trained on 4 NVIDIA A800 GPUs for 1M steps at each stage, with a validation step of 10K. The training pseudocode and the specific prompts used for LLMs are presented in demo page.
@@
5.2 Evaluation Setups
Evaluation Metrics. We evaluate the SDR performance on the commonly used publicly available benchmarks, namely, the in-domain AliMeeting-Eval and AISHELL4-Eval, and the out-of-domain AISHELL5-Eval. Following previous works (alimeeting_2023; diarizationlm_2024; aishell5_2025), we use Character Error Rate (CER), concatenated minimum permutation CER (cpCER), $\Delta\textrm{cp}$, speaker-attributed CER (saCER), and $\Delta\textrm{sa}$ as evaluation metrics.
@@ 5.2 Evaluation Setups
For the SD+ASR paradigm, we use Paraformer-large (paraformer_2022) as the ASR module, which was pre-trained on 60K hours of audio data and achieves SOTA performance on various Mandarin ASR benchmarks. For the SD module, we apply four SOTA open-source SD toolkits, including 3D-Speaker (3d_speaker_2025), Pyannote 3.1 (pyannote_2_2023), Diarizen-base (diarizen_1_2025), and Diarizen-large (diarizen_2_2025). Among the four SD models, 3D-Speaker and Diarizen-large have achieved SOTA performance on different benchmarks, while Pyannote 3.1 and Diarizen-base demonstrate competitive performance. By combining the four SD systems with Paraformer-large, we obtain four SD+ASR baseline systems, namely, 3D-Speaker+Para, Pyannote+Para, Diarizen-base+Para, and Diarizen-large+Para.
SD+ASR+LLM refers to a pipeline where a text LLM is used to correct the outputs from the SD+ASR system. For SD+ASR+LLM baselines, prior works using a text LLM for SD+ASR correction are mainly based on English datasets (diarizationlm_2024; sd_llm_3_2024; sd_llm_2_2025). Following the paradigm, we adapt Qwen2.5-7B-Instruct (qwen2.5_2024) and ChatGPT4.5 (gpt4.5_2025) for correction. We first evaluate their zero-shot performance (z.s.) using prompts following prior work (sd_llm_2_2025), then fine-tune Qwen2.5-7B-Instruct (f.t.) on our training datasets for SD+ASR correction.
5.3 Performance without Speaker Registration
@@
For SD+ASR+LLM baselines pipelined with five models, we incorporate a pre-trained text LLM to correct the mis-assigned speaker labels in the results from the best SD+ASR system (i.e., Diarizen-large+Para). Results show that both Qwen2.5-7B-Instruct and ChatGPT-4.5 perform poorly in zero-shot SDR correction, even degrading the overall performance. This finding is consistent with observations in prior works (sd_llm_3_2024; sd_llm_2_2025). This is mainly due to hallucination of LLMs: despite that we explicitly instruct the LLM to only modify speaker labels, the LLM often alters the content of the speakers’ utterances, leading to a significant increase in CER. Finetuning Qwen2.5-7B-Instruct substantially mitigates the hallucination problem, leading to slight SDR performance improvement over Diarizen-large+Para.
Different from the cascaded architectures of SD+ASR and SD+ASR+LLM, the proposed SpeakerLM is a unified E2E-SDR model that jointly performs speaker diarization and recognition. To explore how much SpeakerLM benefits from the ASR training stage, we first report the CER of SpeakerLM-ASR, which is trained on the ASR data mentioned in Section 3.3. Compared with Paraformer-large, SpeakerLM-ASR is trained on a larger ASR dataset and thus achieves lower CER across all test sets, which naturally contributes to a lower cpCER for SpeakerLM. However, it should be noted that $\Delta\textrm{cp}$ is not affected by the model’s ASR performance, which serves as a reliable metric for evaluating the speaker attribution performance.
@@
Impact of Multi-Stage Training. Figure 4 illustrates the performance of SpeakerLM across the proposed four different training stages (Section 3.3). On both Alimeeting-Eval and Simulation-Test, the performance of SpeakerLM improves progressively, demonstrating the effectiveness of the proposed multi-stage training strategy. However, on AISHELL4-Eval and AISHELL5-Eval, CERs at Stage 2 are higher than those at Stage 1. This is because Stage 2 relies on simulated data, while the simulation process does not incorporate any audio data from these two datasets, leading to a domain mismatch. With four training stages, SpeakerLM demonstrates strong generalizability on the out-of-domain test set, i.e, AISHELL5-Eval. This suggests that the subsequent training stages (Stage 3 and Stage 4), which involve fine-tuning on more realistic and diverse meeting-style data, are crucial for mitigating domain mismatch and enhancing the model’s robustness across different evaluation scenarios. In addition, the ASR performance on simulated data is consistently better than that on real data, which is attributed to the simplicity of the simulated data. Nevertheless, we believe the simulation process remains meaningful, as it closely resembles office or workplace scenarios where a voice assistant is used to record meetings.
@@
5.4 Performance with Speaker Registration